OpenStudy (anonymous):

(x^3-3x^2+6x-18)/(x^2-4x)*(2x-4)/(x^4-3x^3)÷(x^2+2x-8)/(16-x) Ill post a drawing to make it viewable just help me please :)

OpenStudy (anonymous):

\[\frac{ x^3-3x2+6x-18 }{ x^2-4x }*\frac{ 2x-4 }{ x^4-3x^3}\div \frac{ x^2+2x-8 }{ 16-x^2}\]

OpenStudy (anonymous):

I know we gotta factor and stuff first dont know the rest

OpenStudy (anonymous):

haha thanks for the effort yeah ive tried this site but its like Im truly lost steps would've been clearer feel me?

OpenStudy (anonymous):

yea I'm sorry

OpenStudy (anonymous):

not your fault aha but thanks for trying to help :)

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

gimme a minute and ill sort it out for u :)

OpenStudy (anonymous):

just have to simplify right?

OpenStudy (anonymous):

uhm I have to solve

OpenStudy (anonymous):

well in directions says perform the indicatted operations and simplify completely

OpenStudy (anonymous):

yeah basically u just need to do what is indicated and simplify... did u type in the equation correctly?

OpenStudy (anonymous):

yeah I got it thanks I just factored and canceled got it ;) thanks tho for helping and yeah I did besides when I wrote it I forgot to make the last x squared.